# The temporary impact of permanent hiring incentives: Evidence from Italy

### v1

@@ 4.1 Difference in differences
The difference-in-differences (DiD) method is one of the most widely used econometric tools adopted in policy evaluation studies (Callaway and Sant’Anna,, 2021; Wooldridge,, 2023). The main identifying assumption of the DiD is that trends across eligible and not eligible cohorts would have kept evolving in parallel in the absence of the policy (i.e., the parallel trends assumption). However, our repeated cross-sectional setting implies that observations across time periods are not drawn from the same underlying population, imposing an additional no-compositional change assumption (Sant’Anna and Xu,, 2023). This additional assumption entails that samples, though potentially composed of different individuals, should exhibit comparable distributions of observable characteristics. In our case, it seems natural for this condition to hold given that the treatment is only related to the age of the employee, meaning that imbalances of group composition should not be attributed to endogenous worker characteristics. We have nonetheless tested this additional requirement by comparing the distribution of the principal covariates in the annual samples, finding almost no significant compositional differences in observable covariates between groups. The results of this test are available in Appendix A, Table A.1.
@@
where $y$ is the short and longer-term outcome of interest. We detail the analyzed outcomes below in subsection 4.4. The treatment group indicator is denoted by $D_{i}$, indicating age eligibility, while time intercepts are denoted by $R^{k}_{i}$, with $k$ denoting the lags and leads from the 2018 treatment period. The interaction of the treatment with the time variable in 2018 identifies the ITT (Intention to treat) effect $\theta_{1,0}$.
We implement two different specifications for $D_{i}$. The first specification uses a standard binary treatment variable, which applies only to individuals turning 34 or 36, leaving 35-year-olds out of the sample. The variable equals 1 for the eligible age group and 0 for the ineligible one. The second specification follows the approach of Cheng and Hoekstra, (2013), defining the treatment intensity of 35-year-olds (now re-included in the sample) as the share of the year spent at age 34—that is, the proportion of the year in which they were potentially eligible for the incentives. To construct this variable, we account for whether an individual’s birthday occurred before or after the 15th of the month, assigning treatment status at the level of whole months accordingly.
Finally, $X_{i}$ denotes the employee and contract characteristics. Specifically, we include gender, education level, and citizenship for our individual-level variables. Job-level information includes the industrial sector(using the ATECO 1-digit classifications), the skill level (using the Classificazione delle professioni 2011999https://www.istat.it/classificazione/classificazione-delle-professioni/ 1-digit classifications), the place of work (11 levels, including all Tuscany provinces and an out-of-Tuscany class including all the observation with workplace outside of Tuscany), the number of contract’s renewals at the beginning of the year, and the potential duration of the contract (computed as the difference between the end of the observation period and the contract’s start, since the actual duration can not be considered due to its correlation with the outcome).
@@ 4.2 Regression discontinuity design
In terms of comparability with the DiD results, it should be noted that the RD and DiD estimates are not strictly identical, as they capture different treatment effects: the RD measures the local discontinuity in the daily probability of conversion at the eligibility threshold, while the DiD recovers the average annual effect for the eligible cohort. To facilitate comparison, RD estimates can be rescaled (e.g. to yearly probabilities) or expressed relative to the average conversion rate, but they should still be interpreted as local effects at the cutoff.
@@ 4.3 Difference in discontinuities
As a final exercise, we then combine both RD and DiD approaches by comparing conversion rates around the age-cutoff in 2018 against the same effect in the other years in the sample, when the age eligibility threshold was either lower, retroactively implemented, or not applicable altogether. In the presence of additional unobserved policies or treatments that are active at the same age-cutoff of the policy we are interested in (i.e., the hiring incentives) both in 2018 and in the previous years (and under the assumption that the effects of these unobserved treatments are constant in time and do not interact with the effect of hiring incentives in 2018), this setup allows us to wash out these confounding treatments and identify the effects of hiring incentives of 2018.
To do so, we employ a Difference in Discontinuities design (Grembi et al.,, 2016), known also as RD-DiD (henceforth). We use the same sample selection strategy as in the RD model discussed above, but extend the analysis to cover the years 2014–2017.
The specification is the following:
@@ 4.4 Short and longer-term effects
In the evaluation of the immediate take-up of the policy, our primary outcome of interest is a binary variable $y_{i}$ taking one if the job relationship $i$ transitions from temporary to permanent. It is set to 0 in all other cases, such as when the contract remains temporary and ends naturally or when it is terminated by either the employer or the employee. Consequently, the interpretation of the main results focuses on how the policy influences the likelihood of experiencing a conversion compared to any other possible outcome.
For the longer-term medium run effects, it is crucial to examine whether the short-run impacts persist and genuinely translate into stable employment (Desiere and Cockx,, 2022; Saez et al.,, 2021). While in our setting the conversion from fixed-term to permanent contracts might appear to guarantee permanence, it is important to note that permanent contracts can still be terminated. This issue has become even more salient following the introduction of the contratto a tutele crescenti, a new form of open-ended contract introduced in Italy in 2015, which progressively increases employment protection with tenure and allows for easier dismissal in the early years of the relationship. $y_{i}$ denotes a set of a binary variables indicating whether the worker (i) still has the same permanent contract, (ii) still has a permanent contract (no matter where), (iii) still has a contract of any sorts with the same firm or in the same sector and (iv) the number of tays worked $n$ years after the policy. We let $n$ take the values of 1, 2, 3, and 4 years after the observed year. However, given the availability of our data, we had to drop the observations after the policy, since we can not follow the individuals in the fourth year following 2019.
@@ 4.5 Spillover effects
As an initial check for potential spillover effects, we re-estimate the DiD using alternative control groups, replacing the baseline controls with individuals aged 37, 38, or 39 in turn. This allows us to examine whether policy effects are different among those closer to the eligibility threshold, who might be more likely to be substituted by eligible workers. If substitution effects between eligible and non-eligible workers were indeed present, we would expect to observe a significantly smaller policy effect for these additional cohorts, indicating that the people closer to the threshold bore the negative consequences of increased conversions among the eligible.
@@ 5.1.1 Difference in Differences
In columns 4 and 5, we re-include the 35-year olds in the sample and and turn to a continuous treatment denoting the proportion of days in 2018 over which individuals were eligibile to the policy. Here we replicate the first and third formulations of the binary DiD analysis, namely the baseline specification without covariates and the richer specification with covariates and clustered errors. The results we obtain are highly consistent across both definitions of the treatment variable. In particular, the estimated coefficients remain positive, statistically significant, and of comparable magnitude, which reinforces the robustness of the findings and suggests that the evaluation is not overly sensitive to the precise operationalization of the treatment. As in the binary case, the inclusion of covariates strengthens the estimated effect, both in size and statistical significance, further confirming the stability and reliability of the policy’s short-run impact. Furthermore, in the continuous specifications, the effects in 2019, which would have suggested a carryover of the policy impact, virtually disappear.
@@
Table Table 1:
Spillovers
We now turn to studying the substitution between eligible and non-eligible workers. We begin by studying whether our results were sensitive to the choice of the control group. For this purpose, we rely on the standard binary DiD specification, which ensures greater robustness, but introduces alternative controls groups encompassing individuals aged 37 to 39. This approach allows us to test whether non-eligible workers just above the threshold, being the most similar to the eligible group, experienced a differential impact compared with older cohorts. We find that the estimated effects remain very similar, as can be easily noted from the estimates plotted in Figure 4. Parallel trends also generally hold among all cohorts, with the only exception of a small difference in hiring rates between the 34 and 37 years old in 2016. These results suggests the absence of spillover effects, as 36-year-olds are not affected differently from their slightly older peers. The estimated coefficients are shown in Appendix B, in the first three columns of Table B.1.
@@
As a final check, we look at the within-workplace workforce composition of workers of similar age. Specifically, we look at the share of eligible co-workers among the 34 and 36 age cohorts in the same workplace the year before the observation. We then interact this indicator with our treatment variable, following the same approach as in the baseline model, to assess whether conversion rates differed systematically in firms that hired a higher proportion of 34-year-olds compared to those employing slightly older cohorts (i.e., the 36-year-olds). The estimates reported in the last column of Table B.1 show no statistically significant effect, suggesting that the observed increase in conversions was not linked with opportunistic adjustments in hiring practices but rather reflected genuine take-up of the incentive among firms already employing eligible workers.
@@ 5.1.2 Regression Discontinuity and RD-DiD estimation
Table 2 presents the results of a simplified regression discontinuity design on the whole 90-day window. Columns 1 and 2 estimate the model using a linear specification of the running variable, while columns 3 and 4 adopt a quadratic one. Furthermore, in the second and fourth columns, we add the same covariates used for the DiD model. Across all specifications, the estimated effects are positive and statistically significant, and their magnitude remains stable regardless of covariate inclusion. Moreover, the quadratic specification yields results that are highly consistent with the linear one, albeit with slightly reduced statistical significance, likely due to its greater flexibility. Although the estimated coefficients are numerically small, they represent a substantial effect: when expressed relative to the average daily probability of conversion, they correspond to an increase ranging from 67% to 81% around the threshold.
@@
We look again at spillover effects, restricting the sample to firms employing at least two workers turning 35-years old in 2018 (for the RD specification) or in each observation year (for the RD-DiD). For each individual, we construct the spillover variable $S$ as detailed in Section 4 and interact it with the treatment and running variable. Our results are reported in Appendix C, Table C.1.
@@ 5.2 Post-policy effects
Turning to longer-term horizons, Table 6 examines outcomes up to four years after the reform, while analogous estimates for two and three years are provided in the Appendix D, Table D.1 e D.2. Once again, we find no significant policy effects on medium-run employment stability. This suggests that the initial increase in conversions largely failed to consolidate into more persistent gains in terms of open-ended employment relationships.
@@
Standard errors are in parentheses.
@@
Standard errors are in parentheses.
@@ 5.3 Robustness checks
Moving to the regression discontinuity design, to assess the robustness, we conduct a density check of the running variable around the cutoff. Specifically, we employ the rddensity function in R, which allows us to visually inspect the distribution of the running variable and identify any potential discontinuities near the threshold. This approach is complemented by the McCrary test, which provides a formal statistical test for any manipulation of the running variable. The absence of significant jumps in the density at the cutoff indicates that observations are continuously distributed around the threshold, thereby supporting the internal validity of our RDD. The density check is presented in Figure 9, providing an intuitive visual assessment and showing that the data are balanced just above and below the cutoff. Similar figures are also emerge for the entirety of the RD-DiD estimation window in Figure C.1, Appendix C.
@@ 6 Conclusions
Hiring incentives have become a popular policy instrument across modern labor markets, aiming to stimulate open-ended employment through reductions in employer social security contributions. This paper offers a comprehensive analysis of the short and medium-term impact of hiring incentives in Italy, with an eye on heterogeneities and spillover effects.
We exploit a 2018 age eligibility change in the Italian incentive schemes, which temporarily extended eligibility to workers under 35. This setting offers a unique opportunity to evaluate the effectiveness of hiring incentives among workers already employed on temporary contracts, in contrast to most existing studies that focus on workers who are either new or estranged from labor markets. It also isolates incentives targeted at new permanent hires from broader hiring schemes.
@@
Taken together, these findings suggest that while social contributions cuts can generate immediate boosts in permanent hiring, their medium-run effectiveness is limited when applied to workers already in employment relationships. For this group, employer decisions seem to be driven less by structural constraints and more by timing, raising doubts about the cost-effectiveness of temporary expansions in eligibility. Future research should continue to examine not only the conditions under which hiring incentives can create lasting employment gains, but also whether they are capable of doing so at all, particularly in labor markets where temporary contracts represent a structural entry point rather than a transitory stage.
Finally, some limitations should be noted. Tuscany provides a useful case study, with labor market characteristics broadly representative of Italy, but caution is needed when extrapolating results to other contexts, both within Europe and beyond. Moreover, our estimates apply specifically to the 35-year-old cohorts affected by the 2018 reform. While the short-term results are consistent with findings from other settings, the external validity of the medium-run effects remains an open question.

### v2

@@ 4.1 Difference in differences
The DiD method is one of the most widely used econometric tools adopted in policy evaluation studies (Callaway and Sant’Anna,, 2021; Wooldridge,, 2023). The main identifying assumption of the DiD is that trends across eligible and not eligible cohorts would have kept evolving in parallel in the absence of the policy (i.e., the parallel trends assumption). However, our repeated cross-sectional setting implies that observations across time periods are not drawn from the same underlying population, imposing an additional no-compositional change assumption (Sant’Anna and Xu,, 2023). This additional assumption entails that samples, though potentially composed of different individuals, should exhibit comparable distributions of observable characteristics. In our case, it seems natural for this condition to hold given that the treatment is only related to the age of the employee, meaning that imbalances of group composition should not be attributed to endogenous worker characteristics. We have nonetheless tested this additional requirement by comparing the distribution of the principal covariates in the annual samples, finding almost no significant compositional differences in observable covariates between groups. The results of this test are available in Appendix B, Table B.1.
@@
where $y$ is the short and longer-term outcome of interest for individual $i$ observed at time $t$. We detail the analyzed outcomes below in subsection 4.4. The treatment group indicator is denoted by $D_{i}$, indicating age eligibility, while time intercepts are denoted by $R^{k}_{i}$, with $k$ denoting the lags and leads from the 2018 treatment period. The interaction of the treatment with the time variable in 2018 identifies the ITT (Intention to treat) effect $\theta_{1,0}$.
We implement two different specifications for $D_{i}$. The first specification uses a standard binary treatment variable, which applies only to individuals turning 34 or 36, leaving 35-year-olds out of the sample. The variable equals 1 for the eligible age group and 0 for the ineligible one. The second specification follows the approach of Cheng and Hoekstra, (2013), defining the treatment intensity of 35-year-olds (now re-included in the sample) as the share of the year spent at age 34; that is, the proportion of the year in which they were potentially eligible for the incentives. To construct this variable, we account for whether an individual’s birthday occurred before or after the 15th of the month, assigning treatment status at the level of whole months accordingly.
Finally, $X_{i}$ denotes the employee and contract characteristics. Specifically, we include gender, education level (divided between "Elementary school", "Middle school", "High school", "Degree and more", and "No level reported"), and a dummy variable for Italian citizenship for our individual-level variables. Job-level information includes the industrial sector (using the ATECO 1-digit classifications), the skill level (using the Classificazione delle professioni 2011999The Classificazione delle professioni 2011 is the Italian version of the ISCO dataset https://www.istat.it/classificazione/classificazione-delle-professioni/ 1-digit classifications), and the place of work (11 levels, including all Tuscany provinces and an out-of-Tuscany class including all the observations with workplace outside of Tuscany). We also account for the starting month of the contract, to capture potential seasonal or cyclical patterns in the labor market, and the potential duration of the contract, computed as the difference between the end of the year and the contract’s starting date—since the actual duration cannot be used due to its endogeneity with respect to the outcome variable.
@@ 4.2 Regression discontinuity design
In terms of comparability with the DiD results, it is worth noting that the RD and DiD capture different parameters: the RD measures the local discontinuity in the daily probability of conversion at the eligibility threshold, whereas the DiD recovers the average annual effect for the eligible cohort. To facilitate comparison, RD estimates can be rescaled (e.g. to yearly probabilities) or expressed relative to the average conversion rate, but they should still be interpreted as local effects at the cutoff. Accordingly, we expect RD estimates to be larger: RD focuses on the salient, fully treated cutoff day, whereas DiD dilutes the effect by averaging over the year and across workers whose treatment is active only during days of fixed-term employment.
@@ 4.3 Difference in discontinuities
As a final exercise, we then combine both RD and DiD approaches by comparing conversion rates around the age-cutoff in 2018 against the same effect in the other years in the sample, when the age eligibility threshold was either lower, retroactively implemented, or not applicable altogether. In the presence of additional unobserved policies or treatments that are active at the same age-cutoff of the policy we are interested in (i.e., the hiring incentives) both in 2018 and in the previous years (and under the assumption that the effects of these unobserved treatments are constant in time and do not interact with the effect of hiring incentives in 2018), this setup allows us to wash out these confounding treatments and more sharply identify the effects of hiring incentives of 2018.
We then employ a Difference in Discontinuities design (Grembi et al.,, 2016), known also as RD-DiD (henceforth). We use the same sample selection strategy as in the RD model discussed above, but extend the analysis to cover the years 2014–2017. The specification is the following:
@@ 4.4 Short and longer-term effects
In evaluating the immediate take-up of the policy, our primary outcome of interest is a binary variable $y_{i}$, which takes the value of one if the job relationship $i$ transitions from temporary to permanent. It is set to 0 in all other cases, such as when the contract remains temporary and ends naturally or when it is terminated by either the employer or the employee. Consequently, the interpretation of the main results focuses on how the policy influences the likelihood of experiencing a conversion compared to any other possible outcome.
For the longer-term medium run effects, it is crucial to examine whether the short-run impacts persist and genuinely translate into stable employment (Desiere and Cockx,, 2022; Saez et al.,, 2021). While in our setting the conversion from fixed-term to permanent contracts might appear to guarantee permanence, it is important to note that permanent contracts can still be terminated. This issue has become even more salient following the introduction of the "contratto a tutele crescenti", a new form of open-ended contract introduced in Italy in 2015, which progressively increases employment protection with tenure and allows for easier dismissal in the early years of the relationship. $y_{i}$ denotes a set of a binary variables indicating whether the worker (i) still has the same permanent contract, (ii) still has a permanent contract (no matter where), (iii) still has a contract of any sorts with the same firm or in the same sector and (iv) the number of days worked $n$ years after the policy. We let $n$ take the values of 1, 2, 3, and 4 years after the observed year. However, given the availability of our data, we had to drop the observations after the policy, as we are unable to follow the individuals selected in 2019 throughout the subsequent four years.
@@
4.5 Heterogeneous effects
In addition to the overall policy effects, we consider additional specifications to study heterogeneous effects. While some of the interacting dimensions may be endogenous to employment outcomes, identification relies on these characteristics being orthogonal to the implementation of incentives. We consider the following effects:
Macrosector and firm size effects
We first analyze heterogeneous exposure to the policy across industries, examining whether take-up differed by sector, identified by three macrosectors: Agriculture (ATECO section A), Industry (ATECO sections C–F), and Services (ATECO sections G–U), as well as by firm size. Since our data do not provide direct information on firm size, we construct a proxy based on the number of hires recorded within each year.
Temporary contact length effects
Next, we examine heterogeneity by the start date of the temporary contract. In fact, contracts stipulated very close to the end of the year may be less likely to benefit from the incentive, partly because of possible bureaucratic delays, and partly because firms might prefer to convert workers who have spent more time with them. Furthermore, the date of hiring is also correlated with the individual’s exposure to the policy (i.e., the number of days worked under a fixed-term contract).
Demographics and job characteristics
We also analyze heterogeneities by gender and skill level of the occupation, divided into low ("Classificazione delle professioni - CdP"101010”Classificatione delle professioni” is the Italian version of ISCO one digit 8-9), medium (CdP one digit level 4-7), and high (CdP one digit level 1-3).
@@ 4.6 Spillover effects
As an initial check for potential spillover effects, we re-estimate the DiD using alternative control groups, replacing the baseline controls with individuals aged 37, 38, or 39 in turn. This allows us to examine whether estimated policy effects differ when building the counterfactual by using the cohorts closer to the eligibility threshold, who may be more likely to be substituted by eligible workers. If substitution effects between eligible and non-eligible workers were indeed present, we would expect to observe a significantly smaller policy effect when using these additional cohorts as control groups, indicating that individuals closer to the threshold bore the negative consequences of increased conversions among the eligible.
@@ 5.1.1 Difference in Differences
In columns 4 and 5, we re-include the 35-year olds in the sample and and turn to a continuous treatment denoting the proportion of days in 2018 over which individuals were eligibile to the policy. Here we replicate the first and third formulations of the binary DiD analysis, namely the baseline specification without covariates and the richer specification with covariates and clustered errors. The results we obtain are highly consistent across both definitions of the treatment variable. In particular, the estimated coefficients remain positive, statistically significant, and of comparable magnitude, which reinforces the robustness of the findings and suggests that the evaluation is not overly sensitive to the precise operationalization of the treatment. As in the binary case, the inclusion of covariates strengthens the estimated effect, both in size and statistical significance, further confirming the stability and reliability of the policy’s short-run impact. Furthermore, in the continuous specifications, the effects in 2019, which could have suggested an anticipatory effect of the policy threshold reinstatement, virtually disappear.
@@
Spillover effects
We now turn to studying the substitution between eligible and non-eligible workers. We begin by studying whether our results were sensitive to the choice of the control group. For this purpose, we rely on the standard binary DiD specification, which ensures greater robustness while also introducing alternative control groups encompassing individuals aged 37 to 39. This approach allows us to test whether non-eligible workers just above the threshold, being the most similar to the eligible group, experienced a differential impact compared with older cohorts. We find that the estimated effects remain very similar, as can be easily noted from the estimates plotted in Figure 4. Parallel trends also generally hold among all cohorts, with the only exception of a small difference in hiring rates between the 34 and 37-year-olds in 2016. These results suggest the absence of spillover effects, as 36-year-olds are not affected differently from their slightly older peers. The estimated coefficients are shown in Appendix C, in the first three columns of Table C.1.
@@
As a final check, we look at the within-workplace workforce composition of workers of similar age. Specifically, we look at the share of eligible co-workers among the 34 and 36 age cohorts in the same workplace the year before the observation. We then interact this indicator with our treatment variable, following the same approach as in the baseline model, to assess whether conversion rates differed systematically in firms that hired a higher proportion of 34-year-olds compared to those employing slightly older cohorts (i.e., the 36-year-olds). The estimates reported in the last column of Table C.1 show no statistically significant effect, suggesting that the observed increase in conversions was not linked with opportunistic adjustments in hiring practices but rather reflected genuine take-up of the incentive among firms already employing eligible workers.
Heterogeneity
Detailed heterogeneity results are reported in Appendix F.1 and summarized in Figure F.1. As expected, fixed-term contracts in the industrial and service sectors show a higher propensity to be converted into permanent positions compared to those in agriculture (Panel a, Figure F.1). This finding is consistent with the notion that industrial and service firms, which are typically characterized by more stable demand for labor, were more likely or better positioned to take advantage of the subsidy by stabilizing their under-35 employees. Regarding skill intensity, Panel b of the same figure shows that the positive impact on conversions is mainly driven by medium-skill occupations.
To explore the role of firm size, we approximate this dimension by examining the distribution of hirings across years and sectors, using the 75th percentile as a threshold to distinguish between smaller and larger firms. Under this definition, the positive effect of the policy appears to be concentrated primarily among smaller firms, whereas virtually no significant effect is detected for larger firms (Panel c, Figure F.1). A plausible explanation is that larger firms may have perceived the subsidy as insufficiently generous or economically meaningful to alter their hiring behavior.
Contract history also proves to be an important factor. Contracts whose potential duration exceeds the median and are therefore eligible for conversion over a longer period show a significantly higher probability of being transformed into permanent ones (Panel d). This result is consistent with a longer window for firms to screen productivity prior to conversion and with higher effective treatment intensity (i.e., more days on fixed-term contracts, resulting in greater exposure to the incentive).
Finally, when considering gender, the results indicate that the increase in conversion rates is almost entirely driven by male workers (Panel e).
@@ 5.1.2 Regression Discontinuity and RD-DiD estimation
Table 2 presents the results of a simplified regression discontinuity design on the whole 90-day window. Columns 1 and 2 estimate the model using a linear specification of the running variable, while columns 3 and 4 adopt a quadratic one. Furthermore, in the second and fourth columns, we add the same covariates used for the DiD model. Across all specifications, the estimated effects are positive and statistically significant, and their magnitude remains stable regardless of covariate inclusion. Moreover, the quadratic specification yields results that are highly consistent with the linear one, albeit with slightly reduced statistical significance, likely due to its greater flexibility. Although the estimated coefficients are numerically small, they represent a substantial effect: when expressed relative to the average daily probability of conversion, they correspond to an increase ranging from 67% to 81% around the threshold. As expected, the RD estimate of the policy’s effect is higher because it captures the sharp response at the cutoff birthday, whereas the DiD averages outcomes annually across workers who differ in exposure intensity, as treatment applies only on days in fixed-term jobs.
@@
We look again at spillover effects, restricting the sample to firms employing at least two workers turning 35 years old in 2018 (for the RD specification) or in each observation year (for the RD-DiD). For each individual, we construct the spillover variable $S$ as detailed in Section 4 and interact it with the treatment and running variable. Our results are reported in Appendix D, Table D.1.
@@ 5.2 Post-policy effects
Turning to longer-term horizons, Table 6 examines outcomes up to four years after the reform, while analogous estimates for two and three years are provided in the Appendix E, Table E.1 e E.2. Once again, we find no significant policy effects on medium-run employment stability. This suggests that the initial increase in conversions largely failed to consolidate into more persistent gains in terms of open-ended employment relationships.
@@
All the standard errors are clustered at workplace level
@@
All the standard errors are clustered at workplace level
@@
Heterogeneity
We repeat our heterogeneity analysis with respect to the medium-term outcomes to assess whether policy effects re-emerge when mediated by other employers and employee characteristics. However, across all outcomes and time horizons, no clear pattern appers, as the vast majority of estimates show no significant differences. In Figure F.2 and Figure F.3 (Appendix F.1), we report the results of the heterogeneity analysis for one year and four years forward, respectively.
@@ 5.3 Robustness checks
Moving to the regression discontinuity design, to assess the robustness, we conduct a density check of the running variable around the cutoff. Specifically, we employ the rddensity function in R, which allows us to visually inspect the distribution of the running variable and identify any potential discontinuities near the threshold. This approach is complemented by the McCrary test, which provides a formal statistical test for any manipulation of the running variable. The absence of significant jumps in the density at the cutoff indicates that observations are continuously distributed around the threshold, thereby supporting the internal validity of our RDD. The density check is presented in Figure 9, providing an intuitive visual assessment and showing that the data are balanced just above and below the cutoff. Similar figures are also emerge for the entirety of the RD-DiD estimation window in Figure D.1, Appendix D.
@@ 6 Conclusions
This paper provides a comprehensive analysis of the short- and medium-term impacts of hiring incentives for the conversion of fixed-term into open-ended contracts, also considering possible heterogeneity in impacts and spillover effects. We exploit a 2018 age eligibility change in the Italian incentive schemes, which temporarily extended eligibility to workers under 35. This setting offers a unique opportunity to evaluate the effectiveness of hiring incentives among workers already employed on temporary contracts, in contrast to most existing studies that focus on workers who are either new or estranged from labor markets. It also isolates incentives targeted at new permanent hires from broader hiring schemes.
@@
Taken together, these findings suggest that while social contributions cuts can generate immediate boosts in permanent hiring, their medium-run effectiveness is limited when applied to workers already in employment relationships. For this group, employer decisions seem to be driven less by structural constraints and more by timing, raising doubts about the cost-effectiveness of temporary expansions in eligibility. Future research should continue to examine not only the conditions under which hiring incentives can create lasting employment gains, but also whether they are capable of doing so at all, particularly in labor markets where temporary contracts represent a structural entry point.
Acknowledgements
The authors would like to thank the Region of Tuscany for providing access to the data used in this study.
Declaration of Generative AI and AI-Assisted Technologies in the Manuscript Preparation Process
During the preparation of this manuscript, the author(s) made use of AI-assisted technologies to enhance efficiency in editing and grammar checking. Following the use of these tools, the author(s) thoroughly reviewed and revised the content as necessary and assume full responsibility for the final version of the publishedarticle.